# Tube-NeRF: Efficient Imitation Learning of Visuomotor Policies from MPC using Tube-Guided Data Augmentation and NeRFs

### 23 Nov 2023 v1

@@ VI-A Evaluation in simulation
Figure 5: Qualitative evaluation of the trajectory tracking performance of the proposed sensorimotor policy, highlighting the high velocity and 3D motion of the robot, which in turn causes large roll/pitch changes. The eight-shaped trajectory (T1) is the same trajectory used to collect the real-world images and to generate the NeRF employed for Data Augmentation, while the circular trajectory is entirely trained without having collected in the real world a task demonstration. This demonstrates that our approach can be used to train vision-based policies (T2) for which no real-world demonstration has been collected, still achieving low tracking errors (comparable to T1).
Figure 6: Robustness of the policy to visual uncertainties: we show images captured by the onboard camera while using the proposed sensorimotor policy for localization and tracking of T1, with a slung load (tape roll, $0.2$ Kg) attached to the robot. The slung load (circled) repeatedly enters the field of view of the onboard camera, without however compromising the success of the maneuver. (Timestamp added in post-processing).
@@ VI-B Flight experiments
We now validate the data efficiency of Tube-NeRF highlighted in our numerical analysis by evaluating the obtained policies in real-world experiments. We do so by deploying them on an NVIDIA Jetson TX2 (at up to $200$ Hz, TensorRT) on the MIT-ACL multirotor. The policies take as input the fisheye images generated at $30$ or $60$ Hz by the onboard camera. The altitude, velocity and roll/pitch inputs that constitute $\textbf{o}_{\text{other},t}$ are, for simplicity, obtained from the onboard estimator (an EKF fusing IMU with poses from a motion capture system), corrupted with additive noise (zero-mean Gaussian, with parameters as defined in Section V) in the scenarios denoted as “high noise”. We remark that no information on the horizontal position of the UAV is provided to the policy, and horizontal localization must be performed from images. We consider two tasks, tracking the lemniscate trajectory (T1), and tracking a new circular trajectory (denoted T2, velocity up to $2.0$ m/s, duration of $30$ s). No real-world images have been collected for T2, therefore this task is useful to stress-test the novel-view synthesis abilities of the approach, using the NeRF and the nonlinear simulated robot dynamics as a simulation framework. Training. We train one policy for each task, using a single task demonstration collected with DAgger+Tube-NeRF-100 in our NeRF-based simulated environment. During DA, we try to achieve an equal amount between synthetic images (from the NeRF) and real ones (from the database), setting $\bar{\epsilon}=0.5$. Figure 7 reports the number of sampled real images from the database, highlighting that the tube is useful at guiding the selection of real images, but that synthetic images are a key part of the DA strategy (e.g., T2 presents multiple parts without any real image available). Performance under uncertainties. Figure 5 and Table III show the trajectory tracking performance of the learned policy under a variety of real-world uncertainties. Those uncertainties include (i) model errors, such as poorly known drag and thrust to voltage mappings; (ii) wind disturbances, applied via a leaf-blower, (iii) sensing uncertainties (additive Gaussian noise), and (iv) visual uncertainties, produced by attaching a slung-load that repeatedly enters the field of view of the camera, as shown in Figure 6. These results highlight that (a) policies trained after a single demonstration collected in our NeRF-based simulator using Tube-NeRF are robust to a variety of uncertainties while maintaining tracking errors comparable to the ones of the expert (Table III, Figure 2), while reaching velocities up to $3.5$ m/s, and even though the expert localizes using a motion capture system, while the policy uses images from the onboard camera to obtain its horizontal position. In addition, (b) our method enables learning of vision-based policies for which no real-world task demonstration has been collected (e.g., effectively acting as a simulation framework), as shown by the successful tracking of T2, which relied entirely on synthetic training data for large portions of the trajectory (Figure 7), and was obtained using a single demonstration in the NeRF-based simulator. Due to the limited robustness achieved in simulation (Table II), we do not deploy the baselines on the real robot. Efficiency at deployment. Table IV shows that onboard the policy requires on average only 1.5 ms to compute a new action from an image, being at least $\mathbf{5.6}$ faster than a highly-optimized (C/C++) expert. Note that the reported computational cost of the expert is based on the cost of control only (no state estimation), therefore the actual computational cost reduction provided by the policy is even larger.
@@ VII DISCUSSION AND CONCLUSIONS
We have presented Tube-NeRF, a strategy for efficient IL of robust visuomotor policies. Tube-NeRF leverages a robust controller, output feedback RTMPC, to collect demonstrations that account for process and sensing uncertainties. The tube of the controller additionally guides the selection of relevant sensorial observations for DA —considering the effects of sensing noise, domain uncertainties, and disturbances. Extra observations are obtained via a database of real-world data and a NeRF of the environment, while the corresponding actions are efficiently computed using an ancillary controller. We have tailored our approach to localization and control of an aerial robot, showing in numerical evaluations that Tube-NeRF can learn a robust visuomotor policy from a single demonstration, outperforming IL baselines in demonstration and computational efficiency. Our experiments have validated the numerical finding, achieving accurate trajectory tracking using an onboard policy ($1.5$ ms average inference time) that relied entirely on images to infer the horizontal position of the robot, despite challenging 3D motion and uncertainties. In future work, we aim to harness the robustness and efficiency of our policies for localization and control onboard highly agile, computationally constrained, sub-gram aerial robots [40].
VIII ACKNOWLEDGMENTS
Victor D. Li and Tong Zhao for their help with the experiments. Xiaoyi (Jeremy) Cai for the feedback on the manuscript.

### 26 Feb 2024 v2

@@ VI-B Flight experiments
Figure 5: Qualitative evaluation in experiments, highlighting the high velocity and the challenging 3D motion that the student policy can execute under uncertainties.
Figure 6: Robustness to visual uncertainties: we show images captured by the onboard camera while using the proposed sensorimotor policy for localization and tracking of T1, with a slung load (tape roll, $0.2$ Kg) attached to the robot. The slung load (circled) repeatedly enters the field of view of the onboard camera, without however compromising the success of the maneuver. We hypothesize that randomly deleting patches of pixels during training (Sec. IV-B5) contributes to achieving robustness to this disturbance.
@@
We now validate the data efficiency of Tube-NeRF highlighted in our numerical analysis by evaluating the obtained policies in real-world experiments. We do so by deploying them on an NVIDIA Jetson TX2 (at up to $200$ Hz, TensorRT) on the MIT-ACL multirotor. The policies take as input the fisheye images generated at $30$ or $60$ Hz by the onboard camera. The altitude, velocity and roll/pitch inputs that constitute $\textbf{o}_{\text{other},t}$ are, for simplicity, obtained from the onboard estimator (a filter fusing IMU with poses from a motion capture system), corrupted with additive noise (zero-mean Gaussian, with parameters as defined in Section V) in the scenarios denoted as “high noise”. We remark that no information on the horizontal position of the UAV is provided to the policy, and horizontal localization must be performed from images. We consider two tasks, tracking the lemniscate trajectory (T1), and tracking a new circular trajectory (denoted T2, velocity up to $2.0$ m/s, duration of $30$ s). No real-world images have been collected for T2, therefore this task is useful to stress-test the novel-view synthesis abilities of the approach, using the NeRF and the nonlinear simulated robot dynamics as a simulation framework. Training. We train one policy for each task, using a single task demonstration collected with DAgger+Tube-NeRF-100 in our NeRF-based simulated environment. During DA, we try to achieve an equal amount between synthetic images (from the NeRF) and real ones (from the database), setting $\bar{\epsilon}=0.5$. Figure 7 reports the number of sampled real images from the database, highlighting that the tube is useful at guiding the selection of real images, but that synthetic images are a key part of the DA strategy, as T2 presents multiple parts without any real image available. Performance under uncertainties. Figure 5 and Table III show the trajectory tracking performance of the learned policy under a variety of real-world uncertainties. Those uncertainties include (i) model errors, such as poorly known drag and thrust to voltage mappings; (ii) wind disturbances, applied via a leaf-blower, (iii) sensing uncertainties (additive Gaussian noise to the partial state measurements), and (iv) visual uncertainties, produced by attaching a slung-load that repeatedly enters the field of view of the camera, as shown in Figure 6. These results highlight that (a) policies trained after a single demonstration collected in our NeRF-based simulator using Tube-NeRF are robust to a variety of uncertainties while maintaining tracking errors comparable to the ones of the expert (Table III, Figure 2), while reaching velocities up to $3.5$ m/s, and even though the expert localizes using a motion capture system, while the policy uses images from the onboard camera to obtain its horizontal position. In addition, (b) our method enables learning of vision-based policies for which no real-world task demonstration has been collected, effectively acting as a simulation framework, as shown by the successful tracking of T2, which relied entirely on synthetic training data for large portions of the trajectory (Figure 7), and was obtained using a single demonstration in the NeRF-based simulator. Due to the limited robustness achieved in simulation (Table II), we do not deploy the baselines on the real robot. Efficiency at deployment and latency. Table IV shows that onboard the policy requires on average only 1.5 ms to compute a new action from an image, being at least $\mathbf{5.6}$ faster than a highly-optimized (C/C++) expert. Note that the reported computational cost of the expert is based on the cost of control only (no state estimation), therefore the actual computational cost reduction provided by the policy is even larger. Online, image capture (independent of our method, nominal light conditions) has a latency of about $15$ms, while image pre-processing and transfer to the TX2 takes less than $2$ms.. Sensitivity to uncertainties in the visual input. We study the closed-loop performance of the policy under different types of uncertainties in the visual input by monitoring the position tracking error (in simulation, tracking T1, under wind) as a function of different types and magnitudes of noise applied to the images rendered by the NeRF and input to the policy. Specifically, we consider 1) Gaussian noise, representing the presence of high-frequency disturbances/uncertainties such as new visual features in the environment, and 2) Gaussian Blur, capturing visual changes that reduce high-frequency features, such as the ones caused by weather changes, and/or the effects of using a low-quality NeRF for training. The results, shown in Figure 8, highlight that 1) the visual input plays a key role in the output of the policy and, while our approach is not specifically designed for environments with rapidly changing visual appearance, it also shows 2) the overall robustness of the policy to visual uncertainties. Last, the comparison of the effects of the two types of noise highlights that 3) the policy has lower sensitivity to the presence of new high-frequency visual uncertainties, as the position errors grow slower when more Gaussian Noise is applied (lower PSNR) than for Gaussian Blur.
Figure 8: Position tracking error under wind disturbances for T1 in simulation as a function of different types and magnitude of noise in the visual input. PSNR is the Peak Signal to Noise Ratio, and lower PSNR denotes a larger amount of noise corrupting the image. The maximum noise level shown corresponds to to $0\%$ success rate. Contour lines in the Comparison plot show the density of the errors.
@@ VII DISCUSSION AND CONCLUSIONS
We have presented Tube-NeRF, a strategy for efficient IL of robust end-to-end visuomotor policies that achieve real-world robustness in trajectory tracking from images on a multirotors. Tube-NeRF leverages a robust controller, output feedback RTMPC, to collect demonstrations that account for process and sensing uncertainties. In addition, properties of the controller are used to guide a data augmentation procedure, using a combination of a database of real-world images, a NeRF of the environment, and randomization procedures in image space to obtain novel relevant views. For each extra sensorial input, corresponding actions are efficiently computed using an ancillary controller, an integral part of the control framework. We have tailored our approach to localization and control of an aerial robot, showing in numerical evaluations that Tube-NeRF can learn a robust visuomotor policy from a single demonstration, outperforming IL baselines in demonstration and computational efficiency. Our experiments have validated the numerical finding, achieving accurate trajectory tracking using an onboard policy ($1.5$ ms average inference time) that relied entirely on images to infer the horizontal position of the robot, despite challenging 3D motion and uncertainties. In future work, we aim to generalize the approach to dynamic environments by further randomization in image/NeRF space (using [36] for weather conditions), and use event-cameras to ensure performance in poorly-lit environments.